How many gallons of 80% antifreeze solution must be mixed with 100 gallons of 15% antifreeze to get a mixture that is 70% antifreeze?
I need ??? Gallons for answer

Answer by Theo(13342)

let x = the number of gallons of 80% antifreeze.

your formula becomes:

.8 * x + .15 * 100 = .7 * (x + 100)

simplify this formula to get .8 * x + 15 = .7 * x + 70

subtract .7 * x from both sides of the equation and subtract 15 from both sides of the equation to get .8 * x - .7 * x = 70 - 15

simplify to get .1 * x = 55.

solve for x to get x = 55/.1 = 550 gallons.

you would need 550 gallons of 80% antifreeze to mix with 100 gallons of 15% antifreeze to get 650 gallons of 70% antifreeze.

.8 * 550 = 440 gallons of pure antifreeze
.15 * 100 = 15 gallons of pure antifreeze
total = 455 gallons of pure antifreeze

.7 * 650 = 455 gallons of pure antifreeze

solution looks good.

you would need 550 gallons of 80% antifreeze solution.